OpenStudy (goformit100):

2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Solve it?

OpenStudy (goformit100):

@Eyad

OpenStudy (anonymous):

Hmmmm,well with some logic we get 2+3=2*[3+(2-1)]=8 3+7=3*[7+(3-1)]=27 4+5=4*[5+(4-1)]=32 5+8=5*[8+(5-1)]=60 6+7=6*[7+(6-1)]=72 therefore 7+8=7*[8+(7-1)]=98

OpenStudy (goformit100):

yo answer and process seems intresting

OpenStudy (anonymous):

Im pretty sure its 98 o.0